# An Information-Theoretic Regularizer for Lossy Neural Image Compression

### v2

@@ 1 Introduction
We may refer to such a model as the direct coding model for its structural simplicity. Subsequently, following the rate-distortion theorem [1], given a bounded distortion measure $d(\cdot)$, and a distortion constraint $D$, the rate of transmitting $\bm{U}$ is lower bounded by the informational rate-distortion function:
@@
i.e., the minimum mutual information obtained by traversing all the possible encoder-decoder (or quantizer) designs $p_{\bm{\hat{X}}|\bm{X}}$. Theoretically, this bound can be achieved asymptotically using the vector quantization with $N\rightarrow\infty$ [5]. However, as the encoding complexity grows exponentially with the image dimension $N$, a suboptimal yet more practical solution is the transform coding [6]. As depicted in Fig.1(c), for the transform coding model, a deterministic analysis transform $\mathrm{T}_{\mathrm{A}}$: $\mathcal{X}^{N}\rightarrow\mathcal{Y}^{M}$ and a deterministic synthesis transform $\mathrm{T}_{\mathrm{S}}$: $\mathcal{\hat{Y}}^{M}\rightarrow\mathcal{\hat{X}}^{N}$ are inserted into the compression chain. The analysis transform is designed to decorrelate the source $\bm{X}$, such that a much simpler quantization scheme, e.g., uniform scalar quantization: $\bm{u}_{i}=\mathrm{round}(\bm{y}_{i})$, can be independently applied to each dimension of the transformed coefficients $\bm{Y}$. Building on this framework and leveraging the superior nonlinear capabilities of neural networks, numerous neural image compression networks [7, 4, 8, 9, 10, 11, 12, 13] are proposed, continually pushing the frontier of lossy image compression.
Typically, by parameterizing the $\mathrm{T}_{\mathrm{A}}$ and $\mathrm{T}_{\mathrm{S}}$ in Fig.1(c) as neural networks, we have the following rate-distortion training objective [7]:
@@
where the rate $R$ is the estimation of latent entropy $H(\bm{U})$; $q_{\phi}$ is the neural entropy model (or estimated distribution) for $\bm{U}$, whose modeling accuracy can be significantly enhanced by transmitting a small amount of side information $\bm{\hat{Z}}$ (as depicted in Fig.1(c)) as part of the parameters $\phi$ [4]; the distortion $D$ is the $l_{2}$ distance between $\bm{X}$ and $\bm{\hat{X}}$; $\lambda$ is the Lagrangian multiplier, controlling the trade-off of $R$ and $D$. Herein, due to the existence of quantization $\mathrm{Q}(\cdot)$, $\bm{U}$ is discrete in essence. Optimizing the network is then challenging as the gradient is almost zero everywhere for $\mathrm{Q}(\cdot)$. To effectively train the networks, a surrogate is introduced during the training, e.g., additive uniform noise (AUN) [7]:
@@
such that the backward gradient of $\mathrm{Q}(\cdot)$ is one. However, as such approximation inevitably results in the train-test mismatch, and thus high gradient estimation error [14], both the efficiency and efficacy of the optimization are hindered. To address this, many research efforts are devoted to designing advanced surrogates and optimization methods [15, 16, 17, 18, 14].
In this paper, we present a completely distinct view on facilitating the optimization of the compression network: regularization. Conceivably, imposing structural constraints on the network optimization can effectively regularize the gradient estimation error, promote convergence, and reduce overfitting. In the context of generative variational autoencoders [19], structural regularization techniques [20, 21, 22, 23, 24, 25] have been effectively applied to improve authentic data generation. However, since neural compression networks largely differ from generative autoencoders in terms of the minimum latent entropy constraint, to the best of our knowledge, no structural regularization methods tailored to this constraint have been explored. In light of this, in this paper, we propose an information-theoretic regularizer for the neural image compression task. In particular, through an in-depth information-theoretic analysis of compression models, we reveal that minimizing the latent entropy $H(\bm{U})$ is, to some extent, equally important as maximizing the conditional source entropy $H(\bm{X}|\bm{\hat{X}})$. Subsequently, by enforcing the network to simultaneously minimize $-H(\bm{X}|\bm{\hat{X}})$, both the in-domain compact representation and out-of-domain generalization ability are experimentally improved.
@@ 2 Related works
Neural image compression networks are primarily composed of transform, quantization, and latent entropy models. The transform module maps the input source into a latent representation so that the dependencies of the source domain can be more easily handled. Unlike the linear transform [26] used in traditional hybrid coding, the neural transform is inherently nonlinear, allowing for more effective decorrelation of the source data. In the literature, various neural transform architectures are proposed. Representative architectures include the recurrent neural network (RNN) [27], convolutional neural network (CNN) [7], non-local attention [28], transformer [29], mixed transformer-CNN [30], state-space models [13] and invertible networks [31]. Subsequently, the continuous data is quantized into discrete representation, wherein the uniform scalar quantization [7] is commonly adopted. However, as vector quantization in principle can better approach the rate-distortion bound (Eqn.(1)), low-complexity vector quantization [15, 32, 33] and dependent scalar quantization [34, 35] schemes are also studied. Finally, once the quantized latent representations are obtained, a neural entropy model is essential for performing entropy coding. Herein, since the latent can not be perfectly decorrelated, an accurate entropy model that can better capture the dependencies and closely match the ground-truth latent distribution is crucial for the overall network performance. Prominent research directions such as spatial-channel contexts mining [8, 9, 36, 12], expressive latent distribution modeling [10, 37, 38] and latent architecture designs [4, 11, 39] are explored.
@@
The surrogate-based methods [7, 15, 16, 17, 14] prove to be effective in tackling the non-differentiable quantization function during end-to-end optimization. In addition to AUN [7], various quantization surrogates such as soft assignment [15], Gumbel-softmax trick [16], and soft rounding function [17] have been studied. However, as the surrogate introduces a train-test mismatch in forward calculations, the annealing techniques [15, 16, 17, 14] are then proposed to control the “softness” of surrogates, gradually approximating the hard quantization function as training progresses. The straight-through estimator (STE) [40] provides an alternative approach to address the train-test mismatch by enforcing an identity mapping during the backward calculation, thereby enabling direct hard quantization in the forward pass. Nevertheless, STE ultimately introduces a significant gradient estimation bias for the analysis transform. To counter this, an additional fine-tuning stage for the entropy model and synthesis transform has been suggested [18].
While numerous network optimization methods have been developed from a quantization perspective, the regularization view has largely escaped research attention. To the best of our knowledge, the most related efforts are regularization methods for variational autoencoders (VAEs), such as regularization on latent priors [20, 21, 22], reconstruction [23, 24], and consistency [25]. However, as VAEs are fundamentally designed for generative modeling, these methods are not applicable to optimizing compression networks. To bridge this knowledge gap, this paper proposes a structural regularizer for neural image compression and conducts extensive experiments to demonstrate its superiority.
@@ 3.1 Motivation: from minimization to maximization
6:     Compute the backward gradient and update the source entropy model $q_{\theta}$ based on Eqn.(13).
7:end for
@@ 3.2 Information-theoretic regularizer
In principle, optimization neural compression networks with a quantization surrogate, e.g., AUN, introduces gradient estimation errors [14], thereby deviating the objective of minimizing the latent entropy $H(\bm{U})$ under certain distortion constraints (Eqn.(2)), harming both the efficiency and efficacy of the optimization. Motivated by the equivalence of the minimization and maximization optimization discussed in Sec. 3.1, we propose to address this issue by imposing structural maximization constraints on the optimization process. As such, the vanilla objective of minimizing $H(\bm{U})$ is further enhanced through this maximization regularization. Herein, from Theorem 2, the full maximization objective for the transform coding is $H(\bm{X}|\bm{\hat{X}})-H(\bm{U}|\bm{\hat{X}})$, which can be further interpreted as maximizing $H(\bm{X}|\bm{\hat{X}})$ while minimizing $H(\bm{U}|\bm{\hat{X}})$. However, accurately modeling an entropy model $q_{\theta}(\bm{U}|\bm{\hat{X}})$ together with the compression network is quite challenging, as it involves learning another “encoder” function from the reconstruction $\bm{\hat{X}}$ to the latent $\bm{U}$. We experimentally find that a poorly learned $H(\bm{U}|\bm{\hat{X}})$ regularizer can harm the optimization performance, and that the increased training complexity is also burdensome. Consequently, we include only the $H(\bm{X}|\bm{\hat{X}})$ term for regularization, which is an upper bound for the ground-truth $H(\bm{X}|\bm{\hat{X}})-H(\bm{U}|\bm{\hat{X}})$ objective. The entropy modeling is now much easier since the reconstruction $\bm{\hat{X}}$ is closely aligned with the source $\bm{X}$. The resulting training objective is thus given by:
@@
where $R+\lambda D$ follows the rate-distortion loss in Eqn.(2); $-H(\bm{X}|\bm{\hat{X}})$ is the proposed regularizer; $q_{\theta}(\bm{X}|\bm{\hat{X}})$ is the estimated source entropy model; $\alpha$ is regularization factor, controlling the degree of regularization. An illustration for this regularization method is given in Fig.(2). In addition to the compression network, $q_{\theta}(\bm{X}|\bm{\hat{X}})$ is introduced, which is parameterized by a neural network and can be jointly trained via the maximum likelihood criteria [41]:
@@
Overall, the joint training strategy for the compression network and $q_{\theta}(\bm{X}|\bm{\hat{X}})$ is summarized in Algorithm 1, wherein a GAN-style two-stage training strategy [42] is employed.
Figure 3: Performance of the proposed regularization method on three compression models, i.e., hyperprior [4], autoregressive [8], and attention [10]. The anchor is training with the vanilla rate-distortion loss (Eqn.(2)) under an equal number of training steps. $\alpha$ indicates the regularization factor (Eqn.(12)).
@@ 4.1 Experimental setup
Three image compression baseline models are used in the experiments: the hyperprior context model [4] with scale and mean [8], the joint hyperprior and spatial autoregressive context model [8], and the attention architecture [10]. We first train these three models from scratch using the vanilla rate-distortion loss (Eqn.(2)). Then, we compare the results of training these models with the proposed regularization method (Eqn.12) under an equal number of training steps. All three models follow CompressAI’s reimplementation [43], wherein the quantization surrogate is ANU. We may refer to the above three models as hyperprior, autoregressive, and attention, respectively.
@@
For each compression model, we apply exactly the same entropy modeling method as the latent entropy modeling. For instance, for the autoregressive, the $q_{\theta}(\bm{X}|\bm{\hat{X}})$ is also modeled as the Gaussian distribution conditioned on $\bm{\hat{X}}$ with spatial autoregressive context over $\bm{X}$. Our experiments indicate that aligning the entropy model between the latent and source is critical for effective regularization; deviations in either direction (stronger or weaker source models) degrade compression performance. For more details on the network structure of $q_{\theta}(\bm{X}|\bm{\hat{X}})$ and the alignment, readers may refer to the Supplementary Material (Sec. 9 and Sec. 10, respectively).
@@
For training set, we use Flickr20k [44], consisting of 20745 natural images from the Flickr.com. The train-to-validation ratio is set to 9:1. To remove the effect of randomness and facilitate a fair performance comparison, the random seed is fixed to 1 for all training. Following the settings of CompressAI, we train four bit-rate points for each model, i.e., $\lambda\in\{0.0018,0.0035,0.0067,0.0130\}$. The batch size is 16, and the image patch size is $256\times 256$. For optimizer, the experiments use Adam [45], and the learning rates for the compression model and the extra source entropy model are set to $10^{-4}$ and $10^{-3}$, respectively. Herein, following the performance evaluation protocol in prior works [13, 12, 30], we train each model for $2\times 10^{6}$ steps and report the performance, wherein the original attention paper [10] only trains for $1\times 10^{6}$ steps. The performance comparison between the reproduced baselines and CompressAI’s pre-trained models is elaborated in Supplementary Material (Sec. 8). In summary, compared to the pre-trained, our reproduction models yield 0.39%, 0.41%, and -4.11% BD-Rates [46] for the hyperprior, autoregressive, and attention models, respectively, serving as decent reproduction for the experiments. All experiments are conducted on an Nvidia RTX 4090 GPU.
@@
We follow the convention of evaluating over natural image datasets of Kodak [47], CLIC 2024 Validation Set [48] and Tecnick [49]. Besides, to demonstrate generalization performance, we evaluate the models on four out-of-domain datasets: (a) pixel-style dataset [50] (100 images in total); (b) screen-content dataset SCI1K [51] (the first 100 images of the test set); (c) game dataset CCT-CGI [52] (24 images in total); (d) pathology dataset BRACS [53] (the first 100 images of the test set).

### v3

@@ 1 Introduction
We may refer to such a model as the direct coding model for its structural simplicity. Subsequently, following the rate-distortion theorem [1], given a bounded distortion measure $d(\cdot)$ and a distortion constraint $D$, the rate of transmitting $\bm{U}$ is lower bounded by the informational rate-distortion function:
@@
i.e., the minimum mutual information obtained by traversing all the possible encoder-decoder (or quantizer) designs $p_{\bm{\hat{X}}|\bm{X}}$. Theoretically, this bound can be achieved asymptotically using the vector quantization as $N\rightarrow\infty$ [5]. However, since the encoding complexity grows exponentially with the image dimension $N$, a suboptimal yet more practical solution is the transform coding [6]. As depicted in Fig.1(c), for the transform coding model, a deterministic analysis transform $\mathrm{T}_{\mathrm{A}}$: $\mathcal{X}^{N}\rightarrow\mathcal{Y}^{M}$ and a deterministic synthesis transform $\mathrm{T}_{\mathrm{S}}$: $\mathcal{\hat{Y}}^{M}\rightarrow\mathcal{\hat{X}}^{N}$ are incorporated into the compression chain. The analysis transform is designed to decorrelate the source $\bm{X}$, such that a much simpler quantization scheme, e.g., uniform scalar quantization: $\bm{u}_{i}=\mathrm{round}(\bm{y}_{i})$, can be independently applied to each dimension of the transformed coefficients $\bm{Y}$. Building on this framework and leveraging the superior nonlinear capabilities of neural networks, numerous neural image compression networks [7, 4, 8, 9, 10, 11, 12, 13] are proposed, continually pushing the frontier of lossy image compression.
Typically, by parameterizing the $\mathrm{T}_{\mathrm{A}}$ and $\mathrm{T}_{\mathrm{S}}$ in Fig.1(c) as neural networks, the rate-distortion training objective is formulated as [7]:
@@
where $R$ is the estimation of latent entropy $H(\bm{U})$; $q_{\phi}$ is the neural entropy model (or estimated distribution) for $\bm{U}$, whose modeling accuracy can be significantly enhanced by transmitting a small amount of side information $\bm{\hat{Z}}$ (as depicted in Fig.1(c)) as part of the parameters $\phi$ [4]; the distortion $D$ is the $l_{2}$ distance between $\bm{X}$ and $\bm{\hat{X}}$; $\lambda$ is the Lagrangian multiplier, controlling the trade-off between $R$ and $D$. Herein, due to the existence of quantization $\mathrm{Q}(\cdot)$, $\bm{U}$ is discrete in essence. Optimizing the network is then challenging as the gradient is almost zero everywhere for $\mathrm{Q}(\cdot)$. To effectively train the networks, a surrogate is introduced during training, e.g., additive uniform noise (AUN) [7]:
@@
such that the backward gradient of $\mathrm{Q}(\cdot)$ is one. However, since such approximation inevitably results in the train-test mismatch, and thus high gradient estimation error [14], both the efficiency and efficacy of the optimization are hindered. To address this, many research efforts are devoted to designing advanced surrogates and optimization methods [15, 16, 17, 18, 14].
In this paper, we present a completely distinct view on facilitating the optimization of the compression network: regularization. Conceivably, imposing structural constraints on the network optimization can effectively regularize the gradient estimation error, promote convergence, and reduce overfitting. In the context of generative variational autoencoders [19], structural regularization techniques [20, 21, 22, 23, 24, 25] have been effectively applied to improve authentic data generation. However, since neural compression networks fundamentally differ from generative autoencoders in terms of the minimum latent entropy constraint, to the best of our knowledge, no structural regularization methods tailored to this constraint have been explored. In light of this, in this paper, we present the first study of information-theoretic regularization for the neural image compression task. In particular, through an in-depth information-theoretic analysis of compression models, we reveal that minimizing the latent entropy $H(\bm{U})$ is, to some extent, equally important as maximizing the conditional source entropy $H(\bm{X}|\bm{\hat{X}})$. Subsequently, by enforcing the network to simultaneously minimize $-H(\bm{X}|\bm{\hat{X}})$, both the in-domain compact representation and out-of-domain generalization ability are experimentally improved.
@@ 2 Related works
Neural image compression networks are primarily composed of transform, quantization, and latent entropy models. The transform module maps the input source into a latent representation so that the dependencies of the source domain can be more easily handled. Unlike the linear transform [26] used in traditional hybrid coding, the neural transform is inherently nonlinear, enabling more effective decorrelation of the source data. In the literature, various neural transform architectures are proposed. Representative architectures include the recurrent neural network (RNN) [27], convolutional neural network (CNN) [7], non-local attention [28], transformer [29], mixed transformer-CNN [30], state-space models [13] and invertible networks [31]. Subsequently, the continuous transformed coefficients are quantized into the discrete representation, wherein the uniform scalar quantization [7] is commonly adopted. As vector quantization can, in principle, better approach the rate-distortion bound (Eqn.(1)), low-complexity vector quantization [15, 32, 33] and dependent scalar quantization [34, 35] schemes are also studied. Finally, once the quantized latent representations are obtained, a neural entropy model is essential for performing entropy coding. Herein, since the latent cannot be perfectly decorrelated, an accurate entropy model that can better capture the dependencies and closely match the ground-truth latent distribution is crucial for the overall network performance. Prominent research directions such as spatial-channel context mining [8, 9, 36, 12], expressive latent distribution modeling [10, 37, 38] and latent architecture designs [4, 11, 39] are explored.
@@
The surrogate-based methods [7, 15, 16, 17, 14] have been shown to effectively address the non-differentiable quantization function during end-to-end optimization. In addition to AUN [7], various quantization surrogates such as soft assignment [15], Gumbel-softmax trick [16], and soft rounding function [17] have been studied. However, as the surrogate introduces a train-test mismatch in forward calculations, the annealing techniques [15, 16, 17, 14] are then proposed to control the “softness” of surrogates, gradually approximating the hard quantization function as training progresses. The straight-through estimator (STE) [40] provides an alternative approach to address the train-test mismatch by enforcing an identity mapping during the backward calculation, thereby enabling direct hard quantization in the forward pass. Nevertheless, STE ultimately introduces a significant gradient estimation bias for the analysis transform. To counter this, an additional fine-tuning stage for the entropy model and synthesis transform has been suggested [18].
While numerous network optimization methods have been developed from a quantization perspective, the regularization view has largely escaped research attention. To the best of our knowledge, the most related efforts are regularization methods for variational autoencoders (VAEs), such as regularization on latent priors [20, 21, 22], reconstruction [23, 24], and consistency [25]. However, as VAEs are designed for generative modeling, these methods are not applicable to optimizing compression networks. To bridge this knowledge gap, this paper proposes a structural regularizer for neural image compression and conducts extensive experiments to demonstrate its superiority.
@@ 3.1 Motivation: from minimization to maximization
6:     Perform the forward pass of $q_{\theta}$.
7:     Compute the backward gradient and update the source entropy model $q_{\theta}$ based on Eqn.(13).
8:end for
@@ 3.2 Information-theoretic regularizer
In principle, optimizing neural compression networks with a quantization surrogate, e.g., AUN, introduces gradient estimation errors [14], thereby deviating the objective of minimizing the latent entropy $H(\bm{U})$ under certain distortion constraints (Eqn.(2)), harming both the efficiency and efficacy of the optimization. Motivated by the equivalence of the minimization and maximization optimization discussed in Sec. 3.1, we propose to address this issue by imposing structural maximization constraints on the optimization process. As such, the vanilla objective of minimizing $H(\bm{U})$ can be further enhanced through this maximization regularization.
From Theorem 2, an ideal gradient direction for model updating is one that simultaneously maximizes $H(\bm{X}|\bm{\hat{X}})-H(\bm{U}|\bm{\hat{X}})$, which can be further interpreted as maximizing $H(\bm{X}|\bm{\hat{X}})$ while minimizing $H(\bm{U}|\bm{\hat{X}})$. To integrate this regularization into the training objective, two parameterized distributions $q_{\theta}(\bm{X}|\bm{\hat{X}})$ and $q_{\varphi}(\bm{U}|\bm{\hat{X}})$ are necessitated. However, accurately modeling an entropy model $q_{\varphi}(\bm{U}|\bm{\hat{X}})$ alongside the compression network is quite challenging, as it involves learning another “encoder” function from the reconstruction $\bm{\hat{X}}$ to the latent $\bm{U}$. We experimentally find that a poorly learned $H(\bm{U}|\bm{\hat{X}})$ regularizer may harm the optimization performance, and that the increased training complexity is also burdensome. To balance practical feasibility with regularization accuracy, we include only the first $H(\bm{X}|\bm{\hat{X}})$ term for regularization, which is an upper bound for the ground-truth $H(\bm{X}|\bm{\hat{X}})-H(\bm{U}|\bm{\hat{X}})$ objective. The entropy modeling becomes more tractable since the reconstruction $\bm{\hat{X}}$ is closely aligned with the source $\bm{X}$. The primary rationale for such simplification is that although maximizing $H(\bm{X}|\bm{\hat{X}})$ alone is less precise than $H(\bm{X}|\bm{\hat{X}})-H(\bm{U}|\bm{\hat{X}})$, it still provides a meaningful constraint on the ”noisy” gradient, yielding better gradient direction than no regularization. Another justification is that since we will retain the $H(\bm{U})$ minimization term in the overall objective, this $H(\bm{U}|\bm{\hat{X}})$ gap, in principle (Eqn.(11)), should decrease as training progresses, which ensures that the regularization effect remains aligned with the ideal maximization objective. The resulting training objective is thus given by:
@@
where $R+\lambda D$ follows the rate-distortion loss in Eqn.(2); $-H(\bm{X}|\bm{\hat{X}})$ is the proposed regularizer; $q_{\theta}(\bm{X}|\bm{\hat{X}})$ is the estimated source distribution; $\alpha$ is regularization factor, controlling the degree of regularization. An illustration for this regularization method is given in Fig.(2). In addition to the compression network, $q_{\theta}(\bm{X}|\bm{\hat{X}})$ is introduced, which is parameterized by a neural network and can be optimized via the maximum likelihood criteria [41]:
@@
This results in a max-min optimization problem where minimizing Eqn.(12) simultaneously creates a conflict. To resolve this, a GAN-style two-stage joint training strategy [42] is employed: At each training step, we first update the compression network using Eqn.(12), then freeze it and update $q_{\theta}(\bm{X}|\bm{\hat{X}})$ using Eqn.(13). We summarize the strategy in Algorithm 1.
Figure 3: Conditional source entropy modeling for (a) hyperprior [4], (b) autoregressive [8] and attention [10], (c) EILC [43], and (d) MLIC++ [12]. The architecture and module designs are aligned with the latent entropy model. Additional details regarding the specific modules are provided in the Supplementary Material (Sec. 9).
@@ 4.1 Experimental setup
Three classic image compression models are used in exhaustive validation experiments: the hyperprior context model [4] (with scale and mean [8]), the joint hyperprior and spatial autoregressive context model [8], and the attention-based transform model [10]. We may refer to these models as hyperprior, autoregressive, and attention, respectively. Moreover, we conduct the experiments on two advanced compression models: ELIC [43] and MLIC++ [12]. Notably, ELIC utilizes checkerboard and channel-conditioned latent entropy contexts, while MLIC++ builds upon ELIC by introducing additional attention-based context designs and is identified as one of the state-of-the-art (SOTA) models.
We train these models from scratch using the vanilla rate-distortion loss (Eqn.(2)), and compare their performance when trained with the proposed regularization method (Algorithm 1) under an equal number of training steps. For the implementations of the hyperprior, autoregressive, and attention models, we adopt CompressAI’s re-implementation [44], where the quantization surrogate is ANU. For ELIC and MLIC++, we use the open re-implementations from [45] and the official implementation, respectively, both of which employ a mixed ANU-STE quantization surrogate [9].
Figure 4: Performance of the proposed regularization method on the hyperprior [4], autoregressive [8], attention [10], ELIC [43], and MLIC++ [12]. The anchor is trained with the vanilla rate-distortion loss (Eqn.(2)) under an equal number of training steps. $\alpha$ indicates the regularization factor (Eqn.(12)).
@@
For each compression model, we apply exactly the same entropy modeling architecture as the latent entropy modeling $q_{\phi}(\bm{U}|\bm{\hat{Z}})$. Our experiments indicate that aligning the entropy model between the latent and source is critical for effective regularization; deviations in either direction (stronger or weaker source models) degrade compression performance. For the hyperprior model [4], $q_{\theta}(\bm{X}|\bm{\hat{X}})$ is modeled as the factorized Gaussian distribution conditioned on $\bm{\hat{X}}$, i.e.,
$\text{where}\ \bm{\Phi}\triangleq(\mu_{1},...,\mu_{N},\sigma_{1}^{2},...,\sigma_{N}^{2})=f_{\theta}(\bm{\hat{X}}),$ (15)
and $f_{\theta}$ is modeled by convolutional layers; $\mathbb{N}$ denotes the Gaussian distribution; $\bm{\Phi}$ denote the mean and variance; $N$ denotes the dimension of both $\bm{X}$ and $\bm{\hat{X}}$. For both the autoregressive and attention models, the mean and variance are determined by the additional spatial context, i.e.,
$(\mu_{1},...,\mu_{i},\sigma_{1}^{2},...,\sigma_{i}^{2})=f_{\theta}(\bm{\hat{X}},\bm{X}_{<i}),$ (16)
where $\bm{X}_{<i}$ denotes the pixels before the current position and is captured by a $5\times 5$ masked convolution layer. For ELIC, $\bm{\Phi}$ is divided by a checkerboard pattern, leading to the anchor part $\bm{\Phi}_{\text{anchor}}$ and the non-anchor part $\bm{\Phi}_{\text{non-anchor}}$. Both the original frame $\bm{X}$ and $\bm{\Phi}$ are split into 3 slices for context modeling. In the first slice, $\bm{\Phi}_{\text{anchor}}^{1}$ is directly determined by $\bm{\hat{X}}$, while $\bm{\Phi}_{\text{non-anchor}}^{1}$ is determined by an additional spatial context from $\bm{X}^{1}_{\text{anchor}}$. For $k^{th}\ (k>1)$ slice, both the $\bm{\Phi}_{\text{anchor}}^{k}$ and $\bm{\Phi}_{\text{non-anchor}}^{k}$ are conditioned on extra channel context from previous slices $\bm{X}^{<k}$. For MLIC++, following the approach in latent entropy modeling [12], the checkerboard attention module is employed to capture the local dependencies within each slice. Additionally, the intra attention module, which takes the original anchor slice $\bm{X}^{k}_{\text{anchor}}$ and the previous slice $\bm{X}^{k-1}$ as input, captures global dependencies within slices. Furthermore, the inter attention module is introduced to capture the global dependency of inter slices. The detailed network structure of $q_{\theta}(\bm{X}|\bm{\hat{X}})$ for different models is illustrated in Fig. 3. For further details on the convolutional modules, intra-attention module, inter-attention module, and checkerboard attention module, readers may refer to the Supplementary Material (Sec. 9).
@@
For the training set, we use Flickr20k [46], consisting of 20745 natural images from the Flickr.com. The train-to-validation ratio is set to 9:1. To eliminate the effect of randomness and facilitate a fair performance comparison, the random seed is fixed to 1 for all training. Following the settings of CompressAI, we train four bit-rate points for each model, i.e., $\lambda\in\{0.0018,0.0035,0.0067,0.0130\}$. The image patch size is set to $256\times 256$. For the classic hyperprior, autoregressive, and attention models, the batch size is set to 16. The Adam optimizer [47] is employed, with learning rates of $10^{-4}$ for the compression model and $10^{-3}$ for the source entropy model. The training proceeds $2\times 10^{6}$ steps. For the advanced MLIC++ model, the learning rate of the source entropy model is reduced to $10^{-4}$, with a batch size of 8 and training duration of $1\times 10^{6}$ steps. The training setting of the ELIC model is aligned with MLIC++. All experiments are conducted on Nvidia RTX 4090 GPUs.
@@
[…]